Prove or disprove that 2 +(11)^1/2 , beloong to Q
1
Expert's answer
2020-05-15T15:08:59-0400

Let's show that "2 + \sqrt{11} \notin Q" . We will use the fact, that if "r \in Q" , then "r = \frac{p}{q}, q\neq0,p,q \in Z", and "gcd(p,q) = 1" . Let's assume "2 + \sqrt{11} \in Q" . Then "2 + \sqrt{11} = \frac{p_1}{q_1}, q_1 \neq 0, p_1,q_1 \in Z, gcd(p_1, q_1) =1." Hence, "\sqrt{11} = \frac{p_1 - 2q_1}{q_1}." Now, we notice that "q_1" still "\neq 0" , "q_1, p_1 - 2q_1 \in Z" and "gcd(p_1 - 2q_1, q_1) =1". The last equation is true since, using the property of gcd, namely: "gcd(a,b) = gcd(b, a \mod b)" for "a = p_1 - 2q_1, b = q_1," we obtain, that:

"gcd(p_1 - 2q_1, q_1) = gcd(q_1, p_1 \mod q_1) = gcd(p_1, q_1) = 1." Hence, "\sqrt{11} \in Q". Let's now assume that:

"\sqrt{11} = \frac{p}{q}, q \neq 0, p,q \in Z, gcd(p,q) = 1." Then, "11 = \frac{p^2}{q^2}", and "p^2 = 11q^2" . We obtain that "p^2" is divisible by 11, and since 11 is prime, "p" is also divisible by 11. Then "p = 11p_0, p_0 \in Z" . So:"11q^2 = p^2 = 11^2{p_0}^2" , therefore "q^2 = 11{p_0}^2" . So, "q^2" is also divisible by 11, and since 11 is still prime, so is "q" . Hence, "q = 11q_0, q_0 \in Z". But this means that "gcd(p,q) = gcd(11p_0, 11q_0) \ge 11" . So, we obtained the contradiction ("gcd(p,q) = 1") with the assumption "2 + \sqrt{11} \in Q" . Thus, "2 + \sqrt{11} \notin Q" .